### v2

@@ 4.4 Streaming Learning With Language Supervision
Fig. 11 illustrates the streaming learning setup, where language supervision plays a central role Yao (2025). Tasks exhibit temporal coherence, wherein a human and an AI system collaborate toward a shared objective by dynamically allocating control based on their respective strengths. As a result, data arrives sequentially—one sample at a time—and interactions unfold in real time. Each episode begins when the human delegates control to the AI via a user query. The AI then autonomously interacts with tools to attempt task completion. After producing an initial result, the human may intervene to offer feedback—correcting errors and providing guidance. The AI integrates this feedback and continues the task. The episode concludes once the AI successfully completes the task or the human resumes control to begin a new one. In formula, given the user query $\mathbf{q}$, an episode $\tau$ is a sequence of LM’s response $\mathbf{a}_{t}$, tool response $\mathbf{o}_{t}$, and human feedback $f$
$\tau=(\mathbf{q},\mathbf{a}_{0},\mathbf{o}_{0},...,f,...,\mathbf{a}_{t},\mathbf{o}_{t},...,f,...,\mathbf{a}_{T}).$
@@
  • Scalable learning across similar tasks. The system should improve as it accumulates experience. This defines a new scaling axis—performance gains through recurrence. For instance, maintainers resolve GitHub issues 5–18× faster than external contributors Kwa et al. (2025), suggesting the value of task familiarity.
@@
  • Interpretability. Especially in mixed-autonomy settings, the AI’s internal state must remain legible so that humans can anticipate its behavior. Carrying over entire trajectories verbatim may hinder interpretability.
Additionally, in mixed autonomy settings, final accuracy is insufficient—since humans can intervene repeatedly until success. We instead discuss alternative evaluation metrics such as the assistance rate, which is the number of human interventions needed per task. Accordingly, metrics that measure stability and plasticity should reflect the following. Take a task that the AI has already learned. If the assistance rate increases when it sees that task again, it may be forgetting. A stable model should maintain low assistance over time for such tasks. For newly encountered tasks, the model may need help at first, but if the assistance rate drops when similar tasks reappear, it is a sign that the model is learning from its past experience.
@@ 5 Conclusion
This survey introduces a paradigm called training of scaffolded LMs with language supervision. It organizes the intricate literature in prompt optimization, LM pipelines, experiential learning agent, AI workflow optimization, and LM agents. We focus on the structure around LMs where AI and human share control and decision-making responsibilities to achieve a common goal OpenAI (2018); Wang et al. (2024c); Windsurf , and refer to this structure as the scaffold. We view scaffolded LMs as semi-parametric models where the scaffold represents the non-parametric component and the LM refers to the parametric component. From a learning perspective, by focusing on the training of non-parametric variables, we view this new paradigm as learning from language. In practice, scaffolded LM receive instruction, interface with tools, and receive feedback all in natural language. Accordingly, we use an LM as an optimizer to interpret this rich language supervision. In contrast, parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning), using well-defined loss functions. Furthermore, it discusses a key capability missing in current scaffolded LMs: the ability to continuously learn across episodes Silver & Sutton (2025). Parametric training inherently suffers from catastrophic forgetting. Instead, scaffolded LMs are naturally exposed to rich user feedback Wang et al. (2024c); Windsurf . Therefore, we can use an LM optimizer to interpret fuzzy objective Lee et al. (2024) and rich feedback. Compared to parametric training, it is interpretable, efficient, and compatible with closed-source models. Lastly, we hope this survey inspires future research in making humans workflow more efficient through improved mixed autonomy between AI and human.
In addition to learning from language, we can extend learning to incorporate visual information, as some concepts—find me this cup—are difficult to convey purely through text. Multi-modal prompt-based learning has been explored in prior work on robotics Jiang et al. (2023), where prompts may include both text and images, e.g., “bring me image of the cup.” Inspired by this, scaffolded LMs can also interleave textual and visual tokens within their system prompt. This is particularly beneficial when images convey information more compactly or clearly than text alone.
@@
Limitations. Our work focuses on organizing a diverse and intricate body of literature, but it does not aim to exhaustively trace how each idea has been applied across all prior work. For instance, we highlight certain benchmarks that are not consistently used across the papers we discuss Tan et al. (2025). This is because these benchmarks are often too challenging for current non-parametric training methods. This selective approach reflects our intent to extract broader conceptual insights rather than follow any one established framework. With the recent surge of interest in RL applied to LMs, many earlier methods and assumptions are becoming less central. As a result, we intentionally de-emphasize some threads of the literature that, while once influential, no longer align with the direction of current research. This allows us to focus on outlining key future research areas that involves learning from language.

### v3

@@ 5 Beyond episodic learning
Below, we discuss key capabilities required for agents to inhabit streams of experience using non-parametric training with language supervision. To support such learning, we use an LM as an optimizer that interprets rich user feedback, execution traces, and intricate objectives (e.g., “minimize redundant steps based on prior interaction history”, “align with user-specified constraints”), and update non-parametric variables. Compared to parametric training, it uses rich textual feedback and does not suffer from sparse rewards [139] and catastrophic forgetting [3]. Each episode begins when the human delegates control to the AI via a user query. The AI then autonomously interacts with tools to attempt task completion. After producing an initial result, the human may intervene to offer feedback—correcting errors and providing guidance. The AI integrates this feedback and continues the task. The episode concludes once the AI successfully completes the task or the human resumes control to begin a new one. In formula, given the system message and user message $\mathbf{s}_{0}$, an episode $\tau$ is a sequence of LM’s response $\mathbf{a}_{t}$, tool response $\mathbf{o}_{t}$, and human feedback $f$
$\tau=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{o}_{0},...,f,...,\mathbf{a}_{t},\mathbf{o}_{t},...,f,...,\mathbf{a}_{T}).$
@@
  • Scalable learning across similar tasks. The system should improve as it accumulates experience. This defines a new scaling axis—performance gains through recurrence. For instance, maintainers resolve GitHub issues 5-18x faster than external contributors [140], suggesting the value of task familiarity.
@@
  • Interpretability. Especially in mixed-autonomy settings, the AI’s internal state must remain legible so that humans can anticipate its behavior. Carrying over entire trajectories verbatim hinders interpretability.
Figure 12: Continual learning with non-parametric vs. parametric training. Top: we focus on non-parametric training where users provide rich textual feedback at each episode. This feedback is naturally present as users interact with the agent to adjust the answer. We update the agent with an LM optimizer after each episode, allowing agents to learn across episodes [2]. Bottom: parametric training uses scalar feedback and focuses on solving catastrophic forgetting that occurs when the i.i.d assumption doesn’t hold at each gradient step [3].
5.1 Scalability
In mixed-autonomy deployments, users repeatedly face related sub-problems under shared tools, codebases, or data silos. This induces recurrence: the same APIs, preconditions, and failure modes reappear across episodes. If a scaffolded LM fails to improve under recurrence, it repeatedly pays exploration cost and re-commits prior mistakes. We therefore target scaling for tasks that share similar context (e.g., same codebase): systematic gains as similar tasks recur, achieved by updating non-parametric variables (prompts, tools, step policies) with language supervision rather than gradient steps.
Streams of experience naturally provide textual signals—human corrections, rationales, constraint reminders, and execution errors—alongside verifiers (unit tests, checkers, compilers) when available. Such signals (i) localize error causes (wrong tool, argument, or precondition), (ii) support immediate, auditable edits to prompts and tools, and (iii) avoid catastrophic forgetting inherent to weight updates [3].
We highlight two mechanism for scaling in this context:
  • Insight caching (Sec. 4.2.1): distill state-invariant heuristics or decision rules from traces (e.g., “if login required, navigate to /signin before form submit”) and insert them into high-privilege prompts [10, 119].
  • Skill/tool induction (Sec. 4.2.2): refactor recurring action patterns into reusable, typed tools with pre/post-conditions and minimal tests [11, 127, 126, 134, 128].
In-episode reflection [108] improves retries but does not persist if no artifact is updated. RL-tuned reflection controllers (e.g., better verbal feedback and plans) still pay the same exploration cost when a similar task reappears, unless insights or tools are written back [141]. In contrast, agents that inhabit streams of experience synthesize durable artifacts from traces: principle extraction from failures, explicit tool learning.
Prior work equips agents with a memory module [82]. This has long been seen as an important step towards self-evolving agents [142, 143]. For instance, in ExpeL [10], they save successful trajectories in an experience buffer. At inference, it retrieves similar experiences and appends them to the context of the agent. Similarly, Dynamic Cheatsheet [144] stores and retrieves accumulated strategies, code snippets, and general problem-solving insights at inference time. While this helps the agent’s performance, it does not scale well. An important property in the recent breakthrough of AI is the scaling with more data points [145]. We expect the agent to find more efficient solutions as it encounters similar tasks. Retrieving similar experiences is limited to a few experiences. Generative agents [82] use the memory module to simulate real human behaviors that plan and act over very long horizons. In particular, the focus of generative agents is to account for the limited context of LMs by “compressing” previous interactions and retrieving them later. However, these works focus on simply storing experiences [10] or compressing prior interactions [82] to fit multiple interactions beyond the context window.
5.2 Plasticity and Stability
Agents that inhabit streams of experience must adapt quickly (plasticity) while not regressing on previously mastered patterns (stability). Language-level edits (prompts, tools, validators) help by localizing changes, preserving provenance, and enabling rollback—circumventing catastrophic forgetting common in parametric updates [3].
Plasticity and stability requires tracking beyond binary success [146, 147]:
  • Assistance rate: mean human interventions per task. We want to reduce the assistance rate over time without affecting the assistance rate of previous tasks.
  • Backward/forward transfer (BWT/FWT): BWT measures accuracy on prior tasks after an edit; FWT measures improvement on new tasks after learning on previous tasks.
As an analogy, we want to implement new features to a codebase without breaking previous features [50]. We highlight some mechanism to preserve stability:
  • Scoped edits: restrict updates to the smallest locus (one tool, one step prompt) via graph credit assignment [9, 27].
  • Gated promotion: accepte update only if they pass unit tests [50].
  • Versioned artifacts: keep prior versions + provenance (trace, feedback, objective) for audit and instant rollback [11, 126].
5.3 Efficiency
Dumping prior trajectories or long memories into context windows scales poorly as context-windows are limited. Non-parametric training instead compresses experience into durable artifacts—concise insights, executable tools, and step-specific prompts—so future episodes consume similar or fewer tokens while solving harder variants. For instance, we can update prompt and tools with
  • priors: encode stable facts about the environment: API maps, repo layout, naming conventions, data schemas, units, and invariants. This allows fewer exploratory calls.
  • constraints: specifies constraints according to user preferences.
  • docstring: short, copy-ready examples and counter-examples in the tool description.
Besides considering the efficiency of the scaffolded LMs, it is important to also consider the LLM optimizer’s efficiency. Inspired by sleep-time compute [148], we can imagine separating online learning from offline consolidation to preserve latency. In particular, recent work proposed the concept of sleep-time compute [148]. This paper is motivated by the fact that LM user queries consist of a shared task context $c$ and queries. Thus, when the model is idle, it proactively updates the shared context by inferring likely future queries and restructures the context in ways that may be beneficial at test time.
5.4 Interpretability
In mixed autonomy, humans must anticipate and steer agent behavior. Carrying opaque internal states or long, free-form memories across episodes hinders oversight. Non-parametric updates promote legibility: each change is a human-readable diff, on prompt and tools, linked to the trace and feedback that motivated it. In particular, working with language and code allows to easily track:
  • Provenance: instead of storing tensors, it stores pure text.
  • Diffs: relative changes that happen at each step are easily interpretable.
  • Information: the hierarchical structure of the chat format allows to control how information is interpreted [63].
Additionally, several effective approaches operate on the parametric component, offering adaptation but typically with weaker legibility than prompt/tool edits:
Continual learning.
In parametric training, continual learning aims to address a core limitation of parametric training: catastrophic forgetting. Notably, models are optimized using i.i.d. batches sampled from a fixed dataset. However, consider a setting where a model is first trained to classify dogs, and later needs to be extended to also classify cats. Simply continuing training on the new data can lead to the model overwriting its knowledge of dogs—resulting in a phenomenon known as catastrophic forgetting. A naive solution might involve freezing the feature extractor and training only the output head. Yet this often leads to poor performance, as the learned features are not well-suited for the new task (e.g., distinguishing cats). Continual learning addresses this by developing methods that preserve previously learned knowledge while integrating new information, without requiring retraining on the full original dataset. Its motivation is different: to avoid the inefficiency of reprocessing previously seen data—especially when that data has already been well-classified. The key insight is that gradient-based learning on i.i.d. samples can be wasteful when the model only needs to adapt at the margin. Furthermore, gradient-based learning faces the plasticity-stability dilemma [149, 150, 151, 152]: models are highly plastic early in training, quickly adapting to new data, but the plasticity decreases over time. As shown in Fig. 12, besides the aforementioned challenges, continual learning would also require humans to provide labels (in the form of reward or demonstration). In contrast, updating non-parametric variables does not exhibit a decaying plasticity over time and mixed-autonomy settings naturally provides feedback in the form of textual corrections and guidance.
Test-time training.
test-time training methods [153] function as memory mechanisms that compress information directly within the embedding space, rather than relying on text-based representations. For instance, Lattice [153] introduces a compression model that dynamically updates and maintains a compact representation of contextual history by operating on the key-value caches of a transformer in a streaming fashion. These methods typically involve architectural modifications to the neural network and offer an orthogonal approach to standard prompting or finetuning. Potentially, the right objective function for these methods could enable emergent behaviors—such as learning across episodes—in the embedding space.
@@ 6 Conclusion
This survey introduces a paradigm called training of scaffolded LMs with language supervision. It organizes the intricate literature in prompt optimization, LM pipelines, experiential learning agents, AI workflow optimization, and LM agents. We focus on the structure around LMs where AI and human share control and decision-making responsibilities to achieve a common goal [4, 5, 94], and refer to this structure as the scaffold. We view scaffolded LMs as semi-parametric models where the scaffold represents the non-parametric component and the LM refers to the parametric component. From a learning perspective, by focusing on the training of non-parametric variables, we view this new paradigm as learning from language. In practice, scaffolded LMs receive instructions, interface with tools, and receive feedback all in natural language. Accordingly, we use an LM as an optimizer to interpret this rich language supervision. In contrast, parametric training has excelled in learning from demonstration (supervised learning), exploration (reinforcement learning), and observation (unsupervised learning), using well-defined loss functions. Furthermore, it discusses a key capability missing in current scaffolded LMs: the ability to continuously learn across episodes [2]. Parametric training inherently suffers from catastrophic forgetting. Instead, scaffolded LMs are naturally exposed to rich user feedback [5, 94]. Therefore, we can use an LM optimizer to interpret fuzzy objectives [154] and rich feedback. Compared to parametric training, it is interpretable, efficient, and compatible with closed-source models.
We discussed agents that inhabit streams of experience as AI systems that can continuously learn in settings with significant contextual recurrent. From a product perspective [95], LMs are scaffolded to create product that works with the human in a mixed-autonomy settings. We hope this survey inspires future research in adaptive AI systems in mixed-autonomy settings, where AI continously adapts to human feedbacks.
7 Limitations
@@
Our work focuses on organizing a diverse and intricate body of literature, but it does not aim to exhaustively trace how each idea has been applied across all prior work. For instance, we highlight certain benchmarks that are not consistently used across the papers we discuss [155]. This is because these benchmarks are often too challenging for current non-parametric training methods. This selective approach reflects our intent to extract broader conceptual insights rather than follow any one established framework. With the recent surge of interest in RL applied to LMs, many earlier methods and assumptions are becoming less central. As a result, we intentionally de-emphasize some threads of the literature that, while once influential, no longer align with the direction of current research. This allows us to focus on outlining key future research areas that involve learning from language.
In addition to learning from language, there has been limited work in learning to incorporate visual information, as some concepts—find me this cup—are difficult to convey purely through text. Multi-modal prompt-based learning has been explored in prior work on robotics [156], where prompts may include both text and images, e.g., “bring me image of the cup.” Inspired by this, scaffolded LMs can also interleave textual and visual tokens within their system prompt. This is particularly beneficial when images convey information more compactly or clearly than text alone.
[…]
This work was supported by the National Natural Science Foundation of China (Grant Nos. W2442033, W2442032, and 62461160309).